# MC-MKE: A Fine-Grained Multimodal Knowledge Editing Benchmark Emphasizing Modality Consistency

### v1

@@ Abstract
Multimodal large language models (MLLMs) are prone to non-factual or outdated knowledge issues, which can manifest as misreading and misrecognition errors due to the complexity of multimodal knowledge. Previous benchmarks have not systematically analyzed the performance of editing methods in correcting these two error types. To better represent and correct these errors, we decompose multimodal knowledge into its visual and textual components. Different error types correspond to different editing formats, which edits distinct part of the multimodal knowledge. We present MC-MKE, a fine-grained Multimodal Knowledge Editing benchmark emphasizing Modality Consistency. Our benchmark facilitates independent correction of misreading and misrecognition errors by editing the corresponding knowledge component. We evaluate three multimodal knowledge editing methods on MC-MKE, revealing their limitations, particularly in terms of modality consistency. Our work highlights the challenges posed by multimodal knowledge editing and motivates further research in developing effective techniques for this task.
@@ 1 Introduction
With the developments of multimodal large language models (MLLMs) , their application has become widespread across various fields. However, these models struggle with a challenge that the knowledge stored within them could be inaccurate or outdated. This issue manifests in two errors: misreading and misrecognition Cheng et al. (2024). As shown in Figure 1, misrecognition occurs when a model mistakenly identifies an image, such as mistaking Mac Allister as Messi. On the other hand, misreading refers to incorrect textual knowledge, such as misremembering Messi’s football team. Recent researches have introduced knowledge editing in multimodal contexts to address these issues.
@@
Following the conventional definition of knowledge-editing in LLMs, a few studies have proposed benchmarks for knowledge editing in MLLMs Cheng et al. (2024); Huang et al. (2024); Li et al. (2024). However, these benchmarks over-simplify the evaluation of multimodal knowledge editing, and do not distinguish the differences between misreading and misrecognition errorsCheng et al. (2024); Huang et al. (2024). Mixing evaluation of the two types of errors leads to inaccurate assessments of knowledge editing methods in real-world scenarios. Methods may appear to successfully inject objective multimodal knowledge, but actually conduct incorrect edits. Take the misreading error in Figure 1 for an example, where a MLLM misrecognizes the image of Messi to Mac Allister, leading to the erroneous knowledge that "the person in the image plays for Liverpool". If a knowledge editing method falsely injecting an knowledge triplet (Mac Allister, Play for, Inter Miami), it may still achieve great performance on prior benchmarks, since the multimodal knowledge (Image of Messi, Play for, Inter Miami) is actually corrected.
To better handle and evaluate these two types of knowledge editing scenarios, we for the first time define the multimodal knowledge in a decomposed format consist of visual knowledge and textual knowledge. In this way, the misreading and misrecognition errors can be distinguished, and thereby be independently corrected by editing different knowledge components. The decomposition of multimodal knowledge also brings up another requirement Consistency. We believe that a knowledge editing method should always ensure the consistency of knowledge across different modalities. This property is the essential difference between the multimodal knowledge editing and uni-modal knowledge editing.
@@
We evaluate three of most renowned multimodal knowledge editing methods including fine-tuning, MENDMitchell et al. (2022) and IKEZheng et al. (2023) on the three subsets of different editing formats. We find that the performance of these methods is far from satisfaction on MC-MKE. None of them can achieve great performance on all three different editing formats, especially for the consistency metric. It is demonstrated that multimodal knowledge editing is still challenging and requires further exploration.
@@
  • We first propose a decomposed definition of multimodal knowledge according to different multimodal knowledge error types.
@@
  • We conduct experiments with various knowledge editing methods on MC-MKE. The results reveal the limitations of existing methods especially for modality consistency. Different from previous research, we find that editing the corresponding component sometimes yields better performance.
@@ 2.2 Multimodal Models
Multimodal large language models have developed rapidly in recent years. BLIP-2 (Li et al., 2023b) apply Q-Former architecture to transform image input into LLMs input tokens. LLaVA(Liu et al., 2024b) and LLaVA-v1.5(Liu et al., 2024a) utilize linear layers or perceptrons to map the vision features into the inputs of LLMs. Through instruction tuning on BLIP2, InstructBLIP(Dai et al., 2024) gains the ability to follow the instructions on different tasks. Notably, MiniGPT-4Zhu et al. (2023) and MiniGPT-v2Chen et al. (2023) are also powerful LVLMs that exhibit strong performance across various vision-language tasks. There are many other MLLMs such as mPLUG-Owl(Ye et al., 2023), Otter(Li et al., 2023a) and Qwen-VL (Bai et al., 2023). Among all MLLMs, GPT-4V(OpenAI, 2023) is the most powerful one now. We select some of these MLLMs on our research.
@@ 3.2 Definition of MMEdit
We define three different edit formats, IE_edit, SRO_edit, IRO_edit.
@@
Benchmark Edit_formats Edit_requirements
MMEdit
KEBench
MIKE
MC-MKE
Table 1: Comparisons of current multimodal knowledge editing benchmarks, MMEdit (Cheng et al., 2024), KEBench (Wu et al., 2024) and MIKE (Li et al., 2024). IE, SRO, IRO represent different editing formats. ✓ and ✗ mean whether the benchmark can provide data of corresponding editing format. In Fine-grained, ✓ means that the corresponding benchmark is constructed based on fine-grained entity information, while ✗ means that the benchmark is constructed around multimodal task data. Edit_requirements are the properties we expect from a good editing method. ✓ and ✗ indicate whether the benchmark contains the ability to test these properties of editing methods.
SRO_edit SRO_edit is focused on editing specific textual knowledge triplets $(s,r,o)$. When we know the exact way to edit the corresponding textual knowledge tuple $(s,r,o\rightarrow\tilde{o})$, we do not need to find the corresponding multimodal data pair. Instead, we can directly use textual editing way. To ensure consistency in input format of multimodal language models, we use a black image as visual input. Subsequent experiments in appendix A have shown that when using questions generated from textual knowledge as input, the type of input image does not significantly impact the accuracy of the answers. In this case, the model’s textual input is the same as the textual knowledge editing task.
IRO_edit In many multimodal datasets, numerous examples do not present the complete construction information of an instance of multimodal knowledge. We only possess the final multimodal data $(i,r,o)$ and may not be able to accurately decompose it into the corresponding visual knowledge and textual knowledge. Nonetheless, we still need to edit such multimodal knowledge. Even though we may not explicitly identify the corresponding visual knowledge and textual knowledge, an effective method should implicitly understand and update the corresponding knowledge.
Therefore, we hope that a good multimodal knowledge editing method can maintain consistency, even when editing with the final multimodal knowledge input. Theoretically, modifying only $(i,r,o\rightarrow\tilde{o})$ should lead to consistency, whether through $(i,e\rightarrow\tilde{e})$ or $(s,r,o\rightarrow\tilde{o})$. However, there is an issue that there could be many non-unique $e\textquoteright$ ’. Our dataset provides automatically generated reasons to determine it is a modification of $(s,r,o)$. A good editing method should automatically use the provided information to determine that the modification should be implemented on the corresponding textual knowledge triplet in IRO_edit of our benchmark.
@@ 3.3 Requirements of MMEdit Method
Consistency Consistency means that a piece of multimodal knowledge is answered consistently across different modalities after multimodal knowledge editing. In IE_edit, if we modify the corresponding visual knowledge $(i,e\rightarrow\tilde{e})$, consistency means that the corresponding multimodal knowledge should also change as: $(i,\tilde{r},\tilde{o})=(i,e\rightarrow\tilde{e})\times_{\tilde{e}=\tilde{s}}(\tilde{s},\tilde{r},\tilde{o})$. In SRO_edit, if we modify the corresponding textual knowledge $(s,r,o\rightarrow\tilde{o})$ while keeping the visual knowledge unchanged, the corresponding multimodal knowledge will also be modified to $(i,e)\times_{e=s}(s,r,o\rightarrow\tilde{o})=(i,r,o\rightarrow\tilde{o})$. In IRO_edit, due to the reasons mentioned above, our dataset provides information that allows the corresponding multimodal knowledge to change as follows: $(i,r,o\rightarrow\tilde{o})=(i,e)\times_{e=s}(s,r,o\rightarrow\tilde{o})$. The definition of portability is similar to consistency in IE_edit. However, our consistency also includes situation on the SRO_edit and IRO_edit directions.
@@
Generality Generality means that after a piece of multimodal knowledge is edited, the model should not only respond to edited output under the format used for editing. It needs to provide correct edited responses under various generalizations, such as rephrased textual input or different images of the same entity.
@@ 4 MC-MKE Benchmark Construction
Since pure textual knowledge editing datasets are constructed from textual knowledge triplets $(s,r,o)$ and contain editing information $(s,r,o\rightarrow\tilde{o})$, we opt for using the textual knowledge editing dataset MQuAKE as the starting point to construct our multimodal knowledge editing dataset MC-MKE. MQuAKE, as a text knowledge editing dataset, contains knowledge triplets and related editing information. Each instance in MQuAKE corresponds to a textual knowledge triplet and its textual editing information.
@@ 4.1 Data Selection
First, we filter the data using a completely black image paired with generated questions. We selected data that our MLLMs could correctly answer. This step ensures that all the edited knowledge is originally known by the model to make sure we are “editing" instead of “learning". The filtered dataset is referred to as $D_{filter_{1}}$.
Finally, we replaced the subject in the questions generated from $(s,r)$ with “the {category} in the picture” where {category} is the entity type previously generated by ChatGPT, seen in appendix D. If the combined question can be correctly answered by all models, the data is then retained. This step ensures the original multimodal knowledge consistency. The final retained multimodal knowledge $(i,r,o)=(i,e)\times(s,r,o)$ constitutes our knowledge editing source dataset $D_{orig}$.
More details are shown in appendix C.
@@
MiniGPT-v2 FT(Vision) 96.08 2.07 94.52 54.02 10.79
FT(LLM) 95.87 0.78 93.91 93.20 10.80
MEND(Vision) 4.34 26.08 4.23 5.13 6.81
MEND(LLM) 45.21 24.30 44.41 26.17 11.74
IKE 47.50 / 68.76 / 26.41
Table 2: Experimental results on IE_edit data for three editing methods editing two different model components on two MLLMs.
@@ 4.2 Dataset Construction
For a multimodal knowledge $(i,r,o)=(i,e)\times_{e=s}(s,r,o)$ in our filtered multimodal knowledge source dataset $D_{orig}$, we sequentially construct editing data under different editing formats. For IE_edit, our editing inputs consist of images and automaticaly generated questions. We choose to use an entity $\tilde{e}$ of the same category as the entity $e$ as the editing target. For SRO_edit, our editing inputs consist of generated questions, with the editing target being the corresponding new knowledge $\tilde{o}$ given in MQuAKE dataset. We require that $\tilde{o}$ is of the same entity category as $o$. For IRO_edit, our editing input is constructed based on the input from SRO_edit, combined with entity types and templates. The target $\tilde{o}$ is chosen from the corresponding data in the SRO_edit editing dataset. more strict requirements can be seen in appendix C.
Reliability Dataset Construction Our Reliability metric is calculated as shown in the following formula. $D_{e}$ is the editing dataset corresponding to the editing format. For each piece of multimodal knowledge $k=(i,e)\times(s,r,o)$ in $D_{e}$, $\tilde{k}$ is the corresponding edited knowledge. $p_{r}$ is the multimodal input used for testing the Reliability of the corresponding editting format. $t_{r}$ is the target reliability output after knowledge editing. $F$ is the multimodal model, and $\theta_{k\tilde{k}}$ represents the parameters of the model after editing a multimodal knowledge $k\rightarrow\tilde{k}$.
@@
Consistency Dataset Construction Our consistency knowledge editing test data requires constructing according to different editing formats. In IE_edit, consistency is defined as $(i,e\rightarrow\tilde{e})\times_{\tilde{e}=\tilde{s}}(\tilde{s},\tilde{r},\tilde{o})=(i,r,o)\rightarrow(i,\tilde{r},\tilde{o})$. Therefore, we construct the input $p_{c}$ corresponding to the multimodal knowledge $(i,\tilde{r},\tilde{o})$. The edited model should output the corresponding $\tilde{o}$ for this input to ensure consistency. In SRO_edit, we will edit the corresponding textual knowledge triplet $(s,r,o\rightarrow\tilde{o})$, and then construct the input $p_{c}$ for multimodal knowledge $(i,r,\tilde{o})$ based on definition of consistency to test whether the edited model can provide a consistent edited answer $\tilde{o}$. In IRO_edit, for each piece of knowledge $(i,r,o)$, we find its corresponding textual knowledge $(s,r,o)$. After editing the multimodal knowledge $(i,r,o\rightarrow\tilde{o})$, we will analyze whether the corresponding textual knowledge $(s,r,o\rightarrow\tilde{o})$ provides a consistent response.
The consistency score is shown in the following formula. $p_{c}$ is the multimodal input, $\theta_{k\tilde{k}}$ is the edited parameters, $t_{c}$ is the corresponding consistency output in different editing format. Others are the same as (3).
@@
Locality Dataset Construction In the edited datasets for the three editing formats, we used data unrelated to the current editing format but of the same type as locality data. In IE_edit, we randomly selected visual information ($i_{loc}$, $e_{loc}$) different from the current entity in $D_{orig}$ as locality data. In SRO_edit, we randomly selected data $(s_{loc},r_{loc},o_{loc})$ different from the current textual knowledge triplet $(s,r,o)$ in $D_{orig}$ as locality data. In IRO_edit, we randomly selected multimodal knowledge $(i,e)\times_{e=s}(s,r,o)$ where $i,e,s,r$, and $o$ are all different in $D_{orig}$ to form locality data $(i_{loc},e_{loc})\times_{e_{loc}=s_{loc}}(s_{loc},r_{loc},o_{loc})$.
@@
$\text{Score}_{L}=\mathbb{E}_{(k,\tilde{k},p_{l})\sim D_{e}}\left[\mathbbm{1}_{F(p_{l};\theta_{k\tilde{k}})=F(p_{l};\theta)}\right]$ (5)
@@
      MiniGPT-v2       FT(Vision)       82.48       2.36       81.38       1.93
      FT(LLM)       97.34       2.63       96.00       94.49
      MEND(Vision)       4.78       86.53       4.94       6.72
      MEND(LLM)       71.89       19.93       69.79       6.41
      IKE       38.59       58.10       24.78       26.37
Table 3: Experimental results on SRO_edit data for three editing methods editing two different model components on two MLLMs.
Generality Dataset Construction For the three forms of multimodal knowledge editing IE_edit, SRO_edit, and IRO_edit we constructed corresponding generalization test datasets from both image and text perspectives. For the image generalization dataset, we used CLIP to process the images previously crawled from the web. Then, we calculated the relevance between the images and entities using the CLIP model and selected the top 5 most relevant images as the test images for entity image generalization. For the text generalization dataset, we use ChatGPT to rewrite 5 variations of the textual input to serve as the test inputs for text generalization. The prompts can be seen in appendix C.
@@
Edit format IE_edit SRO_edit IRO_edit All
#Data 920 982 982 2884
#Relation 28 30 30 30
#Entity 810 1041 1041 1424
#Alias(avg.) 20.46 17.02 17.02 18.11
#Image 2358 - 1311 2550
#Category 49 76 76 76
Table 4: The statistic of different subsets of MC-MKE. #Entity refers to the total number of entities appeared including $s,o$ and $e$. #Alias refers to the number of answer aliases.
Benchmark statistics Eventually, we create MC-MKE, cosist of a total of 2884 pieces of knowledge across three different edit formats. The associated knowledge involves a large amount of entities and relations, indicating the diversity of MC-MKE. It also has an average of $18.11$ answer aliases per sample, significantly reducing misjudgements of the exact match metrics. More details about dataset statistics are presented in Table 4.

### v2

@@ Abstract
Multimodal large language models (MLLMs) are prone to non-factual or outdated knowledge issues, which can manifest as misreading and misrecognition errors due to the complexity of multimodal knowledge. Previous benchmarks have not systematically analyzed the performance of editing methods in correcting these two error types. To better represent and correct these errors, we decompose multimodal knowledge into its visual and textual components. Different error types correspond to different editing formats, which edit distinct parts of the multimodal knowledge. We present MC-MKE, a fine-grained Multimodal Knowledge Editing benchmark emphasizing Modality Consistency. Our benchmark facilitates independent correction of misreading and misrecognition errors by editing the corresponding knowledge component. We evaluate four multimodal knowledge editing methods on MC-MKE, revealing their limitations, particularly in terms of modality consistency. Our work highlights the challenges posed by multimodal knowledge editing and motivates further research in developing effective techniques for this task.
@@ 1 Introduction
With the developments of multimodal large language models (MLLMs), their application has become widespread across various fields. However, these models struggle with the challenge that the knowledge stored within them could be inaccurate or outdated. This issue manifests in two errors: misreading and misrecognition Cheng et al. (2024). As shown in Figure 1, misrecognition occurs when a model mistakenly identifies an image, such as mistaking Mac Allister as Messi. On the other hand, misreading refers to incorrect textual knowledge, such as misremembering Messi’s football team. Recent researchesCheng et al. (2024) have introduced knowledge editing in multimodal contexts to address these issues.
@@
Following the conventional definition of knowledge-editing in LLMs, a few studies have proposed benchmarks for knowledge editing in MLLMs Cheng et al. (2024); Huang et al. (2024); Li et al. (2024). However, these benchmarks over-simplify the evaluation of multimodal knowledge editing, and do not distinguish the differences between misreading and misrecognition errorsCheng et al. (2024); Huang et al. (2024). Mixing evaluation of the two types of errors leads to inaccurate assessments of knowledge editing methods in real-world scenarios. Methods may appear to inject objective multimodal knowledge successfully but actually conduct incorrect edits. Take the misreading error in Figure 1 for an example, where an MLLM misrecognizes the image of Messi to Mac Allister, leading to the erroneous knowledge that "the person in the image plays for Liverpool". If a knowledge editing method falsely injects a knowledge triplet (Mac Allister, Play for, Inter Miami), it may still achieve great performance on prior benchmarks, since the multimodal knowledge (Image of Messi, Play for, Inter Miami) is actually corrected.
To better handle and evaluate these two types of knowledge editing scenarios, we for the first time define multimodal knowledge in a decomposed format consisting of visual knowledge and textual knowledge in multimodal knowledge editing task. In this way, the misreading and misrecognition errors can be distinguished, and thereby be independently corrected by editing different knowledge components. The decomposition of multimodal knowledge also brings up another requirement Consistency. We believe that a knowledge editing method should always ensure the consistency of knowledge across different modalities. This property is the essential difference between multimodal knowledge editing and uni-modal knowledge editing.
@@
We evaluate four of the most renowned multimodal knowledge editing methods including fine-tuning, MEND Mitchell et al. (2022a), IKE Zheng et al. (2023), and SERAC Mitchell et al. (2022b) on the three subsets of different editing formats. We find that the performance of these methods is far from satisfaction on MC-MKE. None of them can achieve great performance on all three different editing formats, especially for the consistency metric. It is demonstrated that multimodal knowledge editing is still challenging and requires further exploration.
@@
  • We first propose a decomposed definition of multimodal knowledge according to different multimodal knowledge error types in multimodal knowledge editing task.
@@
  • We conduct experiments with various knowledge editing methods on MC-MKE. The results reveal the limitations of existing methods, especially for modality consistency. Different from previous research, we find that editing the corresponding component sometimes yields better performance.
Benchmark Edit_formats Edit_requirements
MMEdit
KEBench
MIKE
MC-MKE
Table 1: Comparisons of current multimodal knowledge editing benchmarks, MMEdit (Cheng et al., 2024), KEBench (Wu et al., 2024) and MIKE (Li et al., 2024). IE, SRO, and IRO represent different editing formats. ✓ and ✗ mean whether the benchmark can provide data of corresponding editing format. In Fine-grained, ✓ means that the corresponding benchmark is constructed based on fine-grained entity information, while ✗ means that the benchmark is constructed around multimodal task data. Edit_requirements are the properties we expect from a good editing method. ✓ and ✗ indicate whether the benchmark contains the ability to test these properties of editing methods.
@@ 2.2 Multimodal Models
Multimodal large language models have developed rapidly in recent years. BLIP-2 (Li et al., 2023b) apply Q-Former architecture to transform image input into LLMs input tokens. LLaVA(Liu et al., 2024b) and LLaVA-v1.5(Liu et al., 2024a) utilize linear layers or perceptrons to map the vision features into the inputs of LLMs. Through instruction tuning on BLIP2, InstructBLIP(Dai et al., 2024) gains the ability to follow the instructions on different tasks. Notably, MiniGPT-4Zhu et al. (2023) and MiniGPT-v2Chen et al. (2023) are also powerful LVLMs that exhibit strong performance across various vision-language tasks. There are many other MLLMs such as mPLUG-Owl(Ye et al., 2023), Otter(Li et al., 2023a) and Qwen-VL (Bai et al., 2023). Among all MLLMs, GPT-4V(OpenAI, 2023) is the most powerful one now. We select some of these MLLMs for our research.
@@ 3.2 Definition of MMEdit
We define three different edit formats, IE_edit, SRO_edit, and IRO_edit.
@@
SRO_edit SRO_edit is focused on editing specific textual knowledge triplets $(s,r,o)$. When we know the exact way to edit the corresponding textual knowledge tuple $(s,r,o\rightarrow\tilde{o})$, we do not need to find the corresponding multimodal data pair. Instead, we can directly use textual editing way. To ensure consistency in the input format of multimodal language models, we use a black image as visual input. Subsequent experiments in appendix A have shown that when using questions generated from textual knowledge as input, the type of input image does not significantly impact the accuracy of the answers. In this case, the model’s textual input is the same as the textual knowledge editing task.
IRO_edit In many multimodal datasets, numerous examples do not present the complete construction information about an instance of multimodal knowledge. We only possess the final multimodal data $(i,r,o)$ and may not be able to accurately decompose it into the corresponding visual knowledge and textual knowledge. Even though we may not explicitly identify the corresponding visual knowledge and textual knowledge, an effective method should implicitly understand and update the corresponding knowledge.
Therefore, we hope that a good multimodal knowledge editing method can maintain consistency, even when editing with the final multimodal knowledge input. Theoretically, modifying only $(i,r,o\rightarrow\tilde{o})$ should lead to consistency, whether through $(i,e\rightarrow\tilde{e})$ or $(s,r,o\rightarrow\tilde{o})$. However, there is an issue that there could be many non-unique $\tilde{e}$. Our dataset provides automatically generated reasons to determine it is a modification of $(s,r,o)$. A good editing method should automatically use the provided information to determine that the modification should be implemented on the corresponding textual knowledge triplet in IRO_edit of our benchmark.
@@ 3.3 Requirements of MMEdit Method
Consistency Consistency means that a piece of multimodal knowledge is answered consistently across different modalities after multimodal knowledge editing as shown in Figure 2. In IE_edit, if we modify the corresponding visual knowledge $(i,e\rightarrow\tilde{e})$, consistency means that the corresponding multimodal knowledge should also change as:
$(i,e\rightarrow\tilde{e})\times_{\tilde{e}=\tilde{s}}(\tilde{s},\tilde{r},\tilde{o})\Rightarrow(i,r,o)\rightarrow(i,\tilde{r},\tilde{o})$
In SRO_edit, if we modify the corresponding textual knowledge $(s,r,o\rightarrow\tilde{o})$ while keeping the visual knowledge unchanged, the corresponding multimodal knowledge will also be modified to:
$(i,e)\times_{e=s}(s,r,o\rightarrow\tilde{o})\Rightarrow(i,r,o)\rightarrow(i,r,\tilde{o})$
In IRO_edit, due to the reasons mentioned above, our dataset provides extra information so that when we edit multimodal knowledge $(i,r,o\rightarrow\tilde{o})$ the corresponding textual knowledge will change as follows:
$(i,r,o\rightarrow\tilde{o})\Rightarrow(i,e)\times_{e=s}(s,r,o)\rightarrow(s,r,\tilde{o})$
@@
Generality Generality means that after a piece of multimodal knowledge is edited, the model should not only output the edited knowledge under the exact input used for editing. It needs to provide correct edited responses under various generalizations, such as rephrased textual input or different images of the same entity.
@@ 4 MC-MKE Benchmark Construction
Since pure textual knowledge editing datasets are constructed from textual knowledge triplets $(s,r,o)$ and contain editing information $(s,r,o\rightarrow\tilde{o})$, we opt for using the textual knowledge editing dataset MQuAKE Zhong et al. (2023) as the starting point to construct our multimodal knowledge editing dataset MC-MKE. MQuAKE, as a text knowledge editing dataset, contains knowledge triplets, related editing information and questions as test input. Each instance in MQuAKE corresponds to a textual knowledge triplet and its textual editing information.
@@ 4.1 Data Selection
We first using a completely black image paired with textual questions in MQuAKE to ask the MLLMs to filter the data to obtain $D_{filter_{1}}$. From $D_{filter_{1}}$, we obtain related images of subjects in the questions from Google, and then apply questions which replace subject with its category to ask MLLMs together with corresponding image to obtain filtered data $D_{filter_{2}}$. Finally, we transform the textual questions into multimodal questions and apply these question to obtain the final multimodal knowledge editing source dataset $D_{orig}$. We utilize some of the filtered-out data to construct a training set for methods requiring training dataset to adjust parameters (such as SERAC). More details about data selection and generation quality assessment can be found in Appendix C.
@@
SERAC 98.48 87.65 68.41 96.41 9.09
MiniGPT-v2 FT(Vision) 98.04 66.43 98.13 91.52 16.67
FT(LLM) 95.76 0.59 93.41 91.48 8.71
MEND(Vision) 7.57 56.73 6.17 5.69 11.36
MEND(LLM) 26.52 67.34 29.19 20.17 4.54
IKE 47.61 / 25.24 / 60.60
SERAC 95.76 83.85 81.48 91.48 8.71
Table 2: Experimental results on IE_edit data for four editing methods editing two different model components on two MLLMs. The highest value is highlighted in bold.
@@ 4.2 Dataset Construction
For multimodal knowledge in our filtered multimodal knowledge source dataset $D_{orig}$, we sequentially construct editing data under different editing formats. For IE_edit, our editing inputs consist of images and automatically generated questions. We choose to use an entity $\tilde{e}$ of the same category as the entity $e$ as the editing target. For SRO_edit, our editing inputs consist of textual questions, with the editing target being the corresponding new knowledge $\tilde{o}$ given in MQuAKE dataset. We require that $\tilde{o}$ is of the same entity category as $o$. For IRO_edit, our editing input is constructed based on the input from SRO_edit, combined with entity types and templates. The target $\tilde{o}$ is chosen from the corresponding data in the SRO_edit editing dataset. more strict requirements can be seen in appendix C.
Reliability Dataset Construction Our Reliability metric is calculated as shown in the following formula. $D_{e}$ is the editing dataset corresponding to the editing format. For each piece of multimodal knowledge $k=(i,e)\times(s,r,o)$ in $D_{e}$, $\tilde{k}$ is the corresponding edited knowledge. $p_{r}$ is the multimodal input used for testing the Reliability of the corresponding editing format. $t_{r}$ is the target reliability output after knowledge editing. $F$ is the multimodal model, and $\theta_{k\tilde{k}}$ represents the parameters of the model after editing a multimodal knowledge $k\rightarrow\tilde{k}$.
@@
Consistency Dataset Construction The construction of our Consistency knowledge editing dataset varies depending on the different editing formats. In IE_edit, consistency is defined as Eq (3). Therefore, after editing visual knowledge, we construct the input $p_{c}$ corresponding to the multimodal knowledge. The edited model should output the corresponding $\tilde{o}$ for this input to ensure consistency. In SRO_edit, consistency is defined as Eq (4). We will edit the corresponding textual knowledge triplet, and then construct the multimodal input $p_{c}$ for multimodal knowledge to test whether the edited model can provide a consistent edited answer $\tilde{o}$ given input $p_{c}$. In IRO_edit, consistency is defined as Eq (5). For each piece of multimodal knowledge, we find its corresponding textual knowledge. After editing the multimodal knowledge, we will analyze whether the corresponding textual knowledge provides a consistent response. The consistency score is shown in the following formula. $p_{c}$ is the multimodal input, $\theta_{k\tilde{k}}$ is the edited parameters, $t_{c}$ is the corresponding consistency output in different editing format. Others are the same as Eq (6).
@@
Locality Dataset Construction In the edited datasets for the three editing formats, we used data unrelated to the current editing knowledge but of the same editing format as locality data. In IE_edit, we randomly selected visual information ($i_{loc}$, $e_{loc}$) different from the current entity in $D_{orig}$ as locality data. In SRO_edit, we randomly selected data $(s_{loc},r_{loc},o_{loc})$ different from the current textual knowledge triplet $(s,r,o)$ in $D_{orig}$ as locality data. In IRO_edit, we randomly selected multimodal knowledge $(i,e)\times_{e=s}(s,r,o)$ where $i,e,s,r$, and $o$ are all different in $D_{orig}$ to form locality data $(i_{loc},e_{loc})\times_{e_{loc}=s_{loc}}(s_{loc},r_{loc},o_{loc})$.
@@
$\text{Score}_{L}=\mathbb{E}_{(k,\tilde{k},p_{l})\sim D_{e}}\left[\mathbbm{1}_{F(p_{l};\theta_{k\tilde{k}})=t_{l}}\right]$ (8)
Edit format IE_edit SRO_edit IRO_edit All
Train #Data 3544 5968 5968 15480
#Relation 37 30 30 37
#Entity 3544 5230 5230 5407
#Alias(avg.) 14.18 13.62 13.62 13.75
#Image 21264 - 20790 22134
#Category 142 342 342 343
#Input Samples 28352 35808 47744 111904
Test #Data 920 982 982 2884
#Relation 28 30 30 30
#Entity 810 1041 1041 1424
#Alias(avg.) 20.46 17.02 17.02 18.11
#Image 2358 - 1311 2550
#Category 49 76 76 76
#Input Samples 15640 11784 16694 44118
Table 3: The statistic of different subsets of MC-MKE. #Entity refers to the total number of entities appeared including $s,o$ and $e$. #Alias refers to the number of answer aliases. #Image in Test refers to the number of filtered images in $D_{orig}$
@@
      SERAC       99.49       89.53       65.05       90.43
      MiniGPT-v2       FT(Vision)       98.78       24.81       97.68       31.67
      FT(LLM)       97.35       2.01       93.73       91.24
      MEND(Vision)       4.37       93.50       3.29       2.74
      MEND(LLM)       2.85       76.96       2.62       3.25
      IKE       30.55       91.26       24.83       21.18
      SERAC       97.35       91.43       75.05       91.24
Table 4: Experimental results on SRO_edit data for four editing methods editing two different model components on two MLLMs. The highest value is highlighted in bold.
Generality Dataset Construction For the three forms of multimodal knowledge editing IE_edit, SRO_edit, and IRO_edit, we constructed corresponding generalization test datasets from both image and text perspectives. For the image generalization dataset, we used CLIP to process the images previously crawled from the web. Then, we calculated the relevance between the images and entities using the CLIP model and selected the top 5 most relevant images as the test images for entity image generalization. For the text generalization dataset, we use ChatGPT to rewrite 5 variations of the textual input to serve as the test inputs for text generalization. The prompts and quality assessment can be seen in appendix C.
@@
Benchmark statistics We create MC-MKE consisting of a training set with 111904 samples and a test set with 44118 samples. Methods such as SERAC can apply training set to adjust their configuration. The test set consists of a total of 2884 pieces of knowledge across three different edit formats. The associated knowledge involves a large number of entities and relations, indicating the diversity of MC-MKE. It also has an average of $18.11$ answer aliases per sample, significantly reducing misjudgments of the exact match metrics. More details about dataset statistics are presented in Table 3.